History of the Islamic Middle East II Final Paper

Dustin Mitchell

March 26, 2001

Chapter 1

Introduction

The famous Sufi poet Jalal al-Din Rumi (1207-1273) was the son of Baha al-Din Valad (b. ca. 1184), who was himself a prominent Sufi preacher and teacher in Balkh. For political reasons the family left Balkh, including Valad’s disciples, in 1218 or 1219. Two years later, the city was crushed by the Mongols. Valad brought his group to settle in Konya, clear across the Dar al-Islam, in 1228, where he quickly assimilated himself into the politics of the region. The Seljuk Sultan, Ala al-Din Kaykobad, surrounded himself with scholars and mystics, so Rumi and his father certainly found fertile ground to continue their spiritual ascent. Valad, however, died in 1231, leaving his position to his son. Already, the political structure of Rum was losing its coherence, and when in 1242 the Mongols were finally making their presence felt in Anatolia, the Konyans had no choice but to pay large sums to the Mongols. In so doing, they also sacrificed their power and control to Ghengis Khan. Although Rumi was far less politically involved than his father had been, he refers repeatedly to the oppressive Mongol presence in his poetry.

Rumi was a prolific writer; his major works are the Divan-i Shams-i Tabriz, his Masnavi, Fihi ma fihi, Majalis-i sab’ah, and Makatib. The Divan is, in general, a collection of scenes and stories from Rumi’s spiritual journey. Rumi attributes many of the poems to Shams-i Tabriz, a fellow Sufi who had a huge effect on Rumi’s life. This effect is, however, beyond the scope of this paper. The Masnavi is a didactic work containing much longer poems, with a total length of over 25,000 verses. It is said to have been started at the suggestion of one of Rumi’s disciples, who then transcribed the entire work over a period of some 13 years [?]. The Fihi ma fihi contains transcriptions of some of Rumi’s lectures, but is not as readable as the Masnavi. The Majalis-i sab’ah is again a collection of lectures, but directed to the general public and thus not assuming knowledge of the Sufi disciplines. The Makatib, finally, is a collection of letters to men of power in Konya, often representing an intervention on Rumi’s part in some conflict between a nobleman and a commoner.

It is interesting to note that Rumi considers lyric poetry and lyric poets to be very lowly: “in my own country and among my own people, there is no occupation more shameful than poetry” [?, 43]. This follows from Sura 26/226, “Poets are followed by erring men. Behold how aimlessly they rove in every valley, preaching what they never practise” [?]. Indeed, he claims in the Fihi that he creates his verses for the entertainment of his friends: “I am affectionate to such a degree that when these friends come to me, for fear that they may be wearied I speak poetry so that they may be occupied with that. Otherwise, what have I to do with poetry? By Allah, I care nothing for poetry, and there is nothing worse in my eyes than that” [?, 5]. Despite this attitude, the Divan, Rumi’s work of lyric poetry, contains over 35,000 verses! Although he claims to write for the entertainment of others, this would seem to be a direct contradiction of the Koran. Perhaps, instead, he uses poetry as the most convenient vehicle for his attempts to describe the indescribable.

This paper will concentrate on the Masnavi and the Divan, as the other works are more difficult to find in English translation. These works reflect Rumi’s spiritual and religious positions quite well, but the Majalis-i sab’ah, and Makatib would be more appropriate sources from which to draw information about Rumi’s relations to his contemporary Konyans. In this paper we use R. A. Nicholson’s translation of the Divan [?] and C. E. Wilson’s translation of the Masnavi [?]. We also make occasional use of translations quoted in the commentaries by William Chittick [?] and Annemarie Schimmel [?, ?]. Citations are be by page in each of these books, since often the secondary references are missing or incomplete.

We are concerned here primarily with Rumi’s theology, and the images through which he expresses this theology. There is a pattern throughout establishing a dichotomy between what I shall call the tangible and the intangible. This difference will be explored through Rumi’s concept of Man and His relation to God. In addition to exploring Rumi’s imagery, we will make some conjectures as to his relations to his contemporaries, Muslim and otherwise; we will also offer some thoughts about the poet’s current popularity.

Man

As a significant part of his teachings, Rumi provides voluminous descriptions of the attributes which characterize a good man. The Divan contains a poem which succinctly summarizes these attributes [?, 29]. The first verse is,

The man of God is drunken without wine, The man of God is full without meat.

The intoxication indicated is an intoxication with spiritual wine. This simile is very common in Rumi’s poetry. Rumi distinguishes three types of intoxicating substances: that derived from grapes, that of sensuality, and that of Love. The former two are to be avoided, as indicated in the Koran, and the latter is to be sought. A verse of the Masnavi states, “Know that every sensual desire is like wine and hashish — it veils the intelligence and stupefies rationality. / Wine is not the only intoxicant of intelligence: Whatever is sensual shuts the eyes and ears” [?, 312]. Both grape-wine and sensuality poison the rational mind, and should thus be avoided. Further, recourse to these wines, which the Koran labels “abominations devised by Satan” [?, 89, 5/90], prevents consumption of the wine of Love: “There are vats of that wine and vats of this. Until you break the first, you will never taste the second!” [?, 313]. The poet is breaking the concept of wine into the tangible type, which is sinful, and the intangible type. He further claims that the two are exclusive. His repeated use of this simile may have been an effort to speak to his contemporaries in images they would understand, while simultaneously urging them to cease their sinful habits.

The man of God is distraught and bewildered,

The man of God has no food or sleep.

Rumi is referring in this verse to the poverty associated with dervish-hood. It is said of Rumi that at age six he “performed extraordinary feats of fasting” [?, xvii]. Regardless of the validity of this story, it is certain that fasting and sleepless vigil-keeping were an integral part of Rumi’s life, and thus in his teaching were a part of the life of the ideal man. To the master, poverty is the lack of material links to the tangible world. Thus dependence on food (including the meat of the first verse) or the self-indulgence of sleep is to be avoided. Rumi claims that once freedom from the mundane is achieved, a man is prepared for the Judgment Day, when the illusory world will disappear: “Of all the different kinds of knowledge, on the day of death only the science of poverty will supply provisions for the way” [?, 187].

The man of God is not of air and earth, The man of God is not of fire and water.

Referring to the fundamental elements composing the tangible world, Rumi here claims that the man of God is not composed of the same substance as other men. This is the nonexistence or intangibility that characterizes a man of God:

How long shall I speak of “below” and “above”? Placelessness is my origin, I am not of place, of I know whence place comes. No, be silent, depart into nonexistence, become naught in nonexistence; behold, how I know things from no-things! [?, 161]

Although a man still exists in this stage, he has done away with his tangible representation – he has neither place nor existence, and casts no shadow: “Annihilation was the adornment of him who said, ’Poverty is my pride.’ Like the flame of a candle, he had no shadow....” [?, 188]. Thus, he no longer has a darkening effect on the vulgar world, but an enlightening one.

The man of God is made wise by the Truth, The man of God is not learned from book.

Rumi divides the intellect into two parts. The lower part is the Partial Intellect, and relates to everything which can be learned from others. However the higher intellect, the Universal Intellect, is closed unto itself and cannot be acquired from any tangible source. It is the Truth, and may only be acquired by a process of gnosis.

The intellect is of two kinds: The first is acquired. you learn it like a boy at school, From books, teachers, reflection and rote, from concepts and from excellent and new sciences. Your intellect becomes greater than that of others, but you are heavily burdened because of your acquisition.... The other intellect is a gift of God. Its fountainhead lies in the midst of the spirit. [?, 35]

This verse creates a very clear division between the tangible knowledge of the Partial Intellect and the intangible, unreachable knowledge of the Science of Religion. Rumi’s attention to this issue may be a reaction to the tendency for faith to become overly intellectual. The passage offers little guidance on becoming wise by the Truth, but it certainly condemns the Partial Intellectual work of his contemporary philosophers and other educated men. Clearly they will be as unprepared as the rest when the judgment day arrives.

Rumi’s outline of the ideal man is hardly appropriate to life in the modern world, but it possesses a very appealing simplicity. Perhaps one cause of the poet’s current popularity is the attraction of so uncomplicated a lifestyle. H. D. Thoureau’s writing appeals to the same sentiment. In particular, both writers’ emphasis on the intangible shows a sharp contrast to the materialism of 20th Century Western culture.

God

Much of Rumi’s verse is spent elucidating and driving home his notions of God. Rumi speaks in apparent contradictions about his God, emphasizing His ineffable nature. Rumi is notably consistent, however, in describing His unity. All things emanate from God, who is the giver of names, and all things will return to him. Rumi’s God is a personal one, in that he communicates with mankind in mankind’s own language. Moreover, when he created the world and named its parts, he gave the names to mankind, giving him power over Creation. Many of his poems describe conversations with his “Beloved”. However, God remains intangible — Rumi bends that distinction slightly through his heavy use of similes, but the division remains. The vulgar world is tangible, and its intangible essence is God. This intangible essence includes the heart of man, which is painfully bound to the tangible world and yearns to be free. Rumi also explores God’s purpose in so binding the hearts of men; it is interesting to reflect on Rumi’s response to the victories of the Mongols in this context.

Rumi makes his use of contradiction explicit in a line from the Divan:

His description is not contained within the intellect, for He is the Coincidence of Opposites. Wonderful Composition without composition! Wonderful freely acting Compelled one! [?, 50]

The human mind cannot comprehend God, because He is composed of opposites. Here, too, we see the unity of God: He is a wonderful Composition, but cannot be represented as the combination of more fundamental parts. Similarly, since He is the Primary Cause for everything, He is certainly a free actor, yet He is the embodiment of Compulsion. One of Rumi’s most common pairs of opposites is that of mercy and wrath: “Such grace and beauty and loveliness and bestowal of life! / O misery and error, if any one dispense with Him!” [?, 117]. God is both the provider of everything good and beautiful, and the punisher of those who have strayed from His path. As noted in the Koran, the earth’s rebirth in the spring is a sign of Divine Grace [?, 83], and His Wrath is evidenced by the misfortunes visited on those on the earth. These misfortunes, however, are for the greater good. In his Divan, Rumi cites Joseph as an example: “Did not Joseph go on a journey from his father, weeping? / Did he not, in the journey, come to fortune and kingdom and victory?” [?, 109]. Indeed, Joseph’s trials proved to be for the good of many people; oftentimes the good result is a spiritual lesson learned personally.

In a pervasive contradiction, Rumi emphasizes the singularity of Reality despite his use of opposites to describe it. In the Divan, he writes:

I have put duality away, I have seen that the two worlds are one;

One I seek, One I know, One I see, One I call.

He is the first, He is the last, He is the outward, He is the inward;^{1}

I know none other except ’Yā Hū’ and ’Yā man Hū [?, 127].

The former two lines present the singularity of God unmistakably, and the latter two again indicate the breadth of His dominion. Rumi likely saw some idol-worship and polytheism among the lower classes in Balkh, and was also probably somewhat offended by Catholicism, the Holy Trinity of which is really a form of tri-theism. Catholicism also has a greater focus on ritual and its associated instruments than does Islam. Often these instruments reflect some part or attribute of God; for instance, communion wafers are physical manifestations of the body of Christ, violating the intangibility of the Divine. To Rumi, all of these forms of worship obscure the true Reality. Thus we find in his Divan:

1The italics indicate a quotation from the Koran: 57/3

I am a painter, a maker of pictures; every moment I shape a beauteous form, And then in thy presence I melt them all away. I call up a hundred phantoms and indue them with a spirit; When I behold thy phantom, I cast them in the fire.

Rumi, fallible like all mankind, constructs images of God or imagines a phantom possessing some part of the Divine, but every time he sets these illusions aside to seek Reality.

God is without place in the world, and man is inextricably tied to his place in the world. However, Rumi does not hesitate to describe intimate conversations between himself and his Lord. God cannot be located:

Seek him in the sky, he shines in water, like the moon; When you come into the water, he will flee to the sky. Seek him in the placeless, he will sign you to place; When you seek him in place, he will flee to the placeless. As the arrow speeds from the bow, like the bird of your imagination, Know that the Absolute will certainly flee from the Imaginary. [?, 81]

If a seeker attempts to find god in some tangible place, he will be greeted only by the tangible illusion — to locate God would be to limit Him, and this is not possible. Man’s heart, however, is tied to the world:

O heart, why art thou a captive in the earth that is passing away?

Fly forth from this enclosure, since thou art a bird of the spiritual world. [?, 175]

The heart belongs outside of place, but has been constrained to the world, creating a fundamental difference between the human soul and God.

The Mongol invasions were generally perceived by Muslims as the Wrath of God wrought on the whole of the dar al-Islam. The Sufis in particular held that governments had strayed far from the boundaries laid out by God, and that whole societies were being punished for their complacency in this error. The souls of thousands are further tied to the tangible world because of their inattention to that which is right. Rumi relates in the Fihi that Ghengis Khan was granted victory because he responded to the execution of his merchants in Khawarazmshah by fasting, weeping, and supplicating himself to God. He was then granted victory by God Himself [?, 33-34]. Although many of the political leaders of Rumi’s era were very pious men, many Sufis claimed that Ghengis Khan, a non-Muslim, was more pleasing to God than the Muslims. A Muslim choosing to follow the path laid out by Rumi, however, would please the Lord and reverse the Mongol tide, and simultaneously free his soul from the tangible world.

Despite the separation of mankind from God, several of the ghazals in the Divan describe personal conversations between Rumi and his Beloved. For example:

He said: ‘Who is at the door?’ Said I: ‘Thy humble slave.’ He said: ‘What business have you?’ Said I: ‘Lord, to greet thee.’ He said: ‘How long will you push?’ Said I: ‘Till thou call.’ He said: ‘How long will you glow?’ Said I: ‘Till resurrection.’ [?, 55]

In this ghazal, Rumi is all but arguing with his Beloved, begging to be allowed to leave his earthly abode and finally be United with God. He shows single-minded purpose, with a ready answer to every question from his Lord. Surprisingly, several verses after this passage Rumi directly contradicts God: “He said: ‘The witness is not valid; your eye is corrupt.’ / Said I: ‘By the majesty of thy justice they are just and clear of sin.”’ This is certainly not a practice of orthodox Islam! Such unorthodoxy is typical of Sufis, and this intimacy may have played a role in the popularity of Sufism: most pagan faiths embody a very intimate relationship between man and the gods. By the time of Rumi’s writing, the Mongols were only partly Islamized. It is likely that pagan Mongols and even first- or second-generation converts felt that the God of Islam was too distant. Rumi’s depiction of God does not bring Him to the level of the tangible, but allows mankind personal contact with the intangible essence of God.

Given the religious trends of the past 50 years, Rumi’s God is a very appealing one to the modern audience. Western culture is progressively usurping the position of God as a controlling figure in the lives of everyday people, and the distant God which has developed among the mainstream monotheistic religions is not adapting to these changes. However, Rumi’s God is in an excellent position to resist this trend: He is singular yet has a part in everything; He gives a purpose to life; He acts directly in the large-scale affairs of mankind; and He is personally approachable. Perhaps it is because these features fit modern circumstances so well that Rumi’s poetry has experienced a modest rebirth in the latter half of this Century.

Conclusion

We have traced a thread of duality through Rumi’s conception of man and his relation to God. The tangible world is physical and inherently limited in every way. God is the intangible, which can neither be understood nor perceived. A good man is one who sheds his tethers to the tangible world and ascends to Unity with God:

Thee I chose, of all the world, alone; Wilt thou suffer me to sit in grief? My heart is as a pen in thy hand, Thou art the cause if I am glad or melancholy. [?, 121]

Rumi desires to become a servant to God; he would reflect the every desire of his Beloved. To do this, the good man must shed his own earthly desires, and find release from all the knowledge he has acquired, which is as tangible as the earth and is of no use in the service of God. He must also develop his Universal Intellect, so that he might understand the Truth and to know how to achieve his union with God.

Rumi’s God, consummately singular despite His description in contradictions and opposites, is completely intangible. Even so, he is a personal God, more approachable than the orthodox God. He created the tangible world by naming it, and made man the master of this world by teaching the names to mankind. “He taught Adam the names of all things” [?, 13, 2/31]. In this sense, Islam allows mankind to manipulate the tangible world, and indeed justifies these manipulations in the name of the Almighty. In his writings relevant to this issue, Rumi emphasizes the need to do what is pleasing to God. Mankind has been given power over the universe, but is expected to use that power wisely.

Among those in the modern West who have foregone religion, anger is often voiced over the ends which mainstream religions have justified throughout history. Such individuals generally frown upon those wage religious wars, advocate colonization, or proselytize. Rumi’s recent popularity may be due in part to his emphasis on personal responsibility and personal devotion.